Yellow:

The sum of three consecutive numbers is 57. What is the largest of these numbers? A) 18 B) 19 C) 20 D) 21

mw104969:

Let the 3 consecutive odd numbers be x, x+2,x+4 x+x+2+x+4=57

mw104969:

3x=51 and from there you can get the answer

TETSXPREME:

How Come?

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME How Come? \(\color{#0cbb34}{\text{End of Quote}}\) if you think thats right then what are the other numbers??

TETSXPREME:

There is Only one answer c:

Yellow:

Oh dangm my bad

mw104969:

ik fore sure its 21 i did this before and got it right

mw104969:

for sure*

mw104969:

What are the answer choices....if The answer choices are 21,19, and 17.............21 is the answer

Yellow:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 What are the answer choices....if The answer choices are 21,19, and 17.............21 is the answer \(\color{#0cbb34}{\text{End of Quote}}\) its 18, 19, 20 and 21 though

TETSXPREME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 What are the answer choices....if The answer choices are 21,19, and 17.............21 is the answer \(\color{#0cbb34}{\text{End of Quote}}\) 17 isnt once of the answer choices

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yellow \(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 What are the answer choices....if The answer choices are 21,19, and 17.............21 is the answer \(\color{#0cbb34}{\text{End of Quote}}\) its 18, 19, 20 and 21 though \(\color{#0cbb34}{\text{End of Quote}}\) im 100% sure its 21 do the math yourself with the steps I showed you if you dont believe me

jhonyy9:

The sum of three consecutive numbers is 57. x +(x+1)+(x+2) = 57 solve this for x and get the x+2 hope helped

mw104969:

im saying 17,19, and 21 would equal to 57

jhonyy9:

CONSECUTIVE NUMBERS - WHAT MEAN ?

Yellow:

My head is blown

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 im saying 17,19, and 21 would equal to 57 \(\color{#0cbb34}{\text{End of Quote}}\) THESE ARE CONSECUTIVE ?

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yellow My head is blown \(\color{#0cbb34}{\text{End of Quote}}\) if you want the points, Im telling you to pick 21 but ayy its your choice so you can pick what you want

Yellow:

Ok wait give me a second

TETSXPREME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Yellow My head is blown \(\color{#0cbb34}{\text{End of Quote}}\) if you want the points, Im telling you to pick 21 but ayy its your choice so you can pick what you want \(\color{#0cbb34}{\text{End of Quote}}\) 21 is Incorrect XDD

Yellow:

It was 20 lolll

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME \(\color{#0cbb34}{\text{Originally Posted by}}\) @mw104969 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Yellow My head is blown \(\color{#0cbb34}{\text{End of Quote}}\) if you want the points, Im telling you to pick 21 but ayy its your choice so you can pick what you want \(\color{#0cbb34}{\text{End of Quote}}\) 21 is Incorrect XDD \(\color{#0cbb34}{\text{End of Quote}}\) how come??

Yellow:

It's alright, I got a good grade

Yellow:

Thank you jhony, mw, and TETSXPREME.

TETSXPREME:

No Porblemo

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yellow Thank you jhony, mw, and TETSXPREME. \(\color{#0cbb34}{\text{End of Quote}}\) you are welcome

jhonyy9:

NP ANYTIME just divide 57/3 = 19 +/-1 = 18 +19 +20 = 57

Yellow:

Ohh okay thank you so very much, makes sense :-)

jhonyy9:

@Laylalyssa clearly ?

Laylalyssa:

thats it/.?

jhonyy9:

yeah - this is so easy

Laylalyssa:

oh ya good job

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa oh ya good job \(\color{#0cbb34}{\text{End of Quote}}\) ty

Laylalyssa:

yw